The Henry's Law constant of methyl bromide, CH3Br, is k = 0.159 mol/(L • atm) at 25°C. What is the solubility of methyl bromide in water at 25°C and at a partial pressure of 270. mm Hg?
Answer #1

Partial pressure of the \(\mathrm{CH}_{3} \mathrm{Br}, P_{\mathrm{cH}_{3} \mathrm{Er}}=270 \mathrm{mmHg} \times \frac{1 \mathrm{~atm}}{760 \mathrm{mmHg}}=0.3553 \mathrm{~atm}\)

Henry's constant of \(\mathrm{CH}_{3} \mathrm{Br}, k_{\mathrm{h}}=0.159 \mathrm{~mol} / \mathrm{L}\). atm

Molar solubility of \(\mathrm{CH}_{3} \mathrm{Br}, C_{0_{2}}=k_{\mathrm{h}} P_{\mathrm{cH}_{3} \mathrm{Er}}\)

$$ \begin{array}{l} =\left(0.159 \frac{\mathrm{mol}}{\mathrm{L} \cdot \mathrm{atm}}\right)(0.3553 \mathrm{~atm}) \\ =0.0565 \mathrm{~mol} / \mathrm{L} \end{array} $$